Ellen's age is half her mother's age
Ellen's father is 5 years older than her mother
Total sum of their ages is 100
What is age of :
Ellen
Her mother
Her father

Answer by Mini-(17)

Let Ellen is X years old, then her mother is 2*X years old,
and her father is 2*X+5 years old.
Total sum of their ages is 100 = X + 2*X + 2*X + 5.
100 = 5*X + 5. 5*X = 100 - 5 = 95.
X = 95/5 = 19.
Ellen's age is 19, her mother is 38 years old and her father is 43 years old.